OpenStudy (anonymous):

A function f is known to have an inverse f^-1. What could be the graph of y =f(x) ?

OpenStudy (campbell_st):

find the graph that is reflected in the line y = x

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

What campbell said.

OpenStudy (anonymous):

How to do that?

OpenStudy (anonymous):

What does that mean

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

To be honest, it's kinda difficult to explain.

OpenStudy (campbell_st):

ok.... on the graph... put a ruler along or draw the line y = x so it goes through (0,0), (1, 1), (2, 2)... find the line that is a mirror image... and thats the inverse function

OpenStudy (anonymous):

And go to number 13

OpenStudy (ipwnbunnies):

I can't get in there because I'm not a Texas student lol.

OpenStudy (campbell_st):

so when you have the line.... find a pair of graphs that are mirror images of each other...

OpenStudy (campbell_st):

so if you folded the graph along the line y = x, which pair of graphs sit exactly on top of each other... and thats the answer

OpenStudy (anonymous):

There are two problems like that I mean answers

OpenStudy (ipwnbunnies):

Yeah, graphically, you can do what campbell is saying. For problems, I believe you exchange x and y, and solve for y.

OpenStudy (anonymous):

Campbell can I send a picture to your email

OpenStudy (anonymous):

@campbell_st

OpenStudy (ipwnbunnies):

For example, find the inverse of f(x) = x^2. y = x^2 -> x = y^2 -> y = +/- sqrt(x) f^-1(x) = +/- sqrt(x)

OpenStudy (anonymous):

Can I send a picture to your email @iPwnBunnies

OpenStudy (ipwnbunnies):

Just attach the picture here.

OpenStudy (anonymous):

I can't I'm on my iPad